# Bridging Privacy and Robustness for Trustworthy Machine Learning

### v4

@@ II-E Relation between Robustness and Privacy
    when corrupted ratio $\tau\approx\log(n)/n\epsilon$, private estimator is naturally robust, robust estimator can be converted to private estimator.
@@
    when corrupted ratio $\tau\geq\frac{2(d\cdot\log(\frac{R}{\alpha_{0}}+1)+\log\frac{2}{\beta})}{n\epsilon}$, robust estimator can be converted to nearly private estimator.
III Privacy Attack and Defense Framework
This section delineates the adversarial landscape confronting data privacy in machine learning applications. We systematically outline the threat model, characterizing the adversary’s objectives, capabilities, knowledge base, and methodologies for launching attacks. Concurrently, we describe the defender’s strategies that underpin the protection mechanisms aimed at safeguarding sensitive information. It is worth noting that this privacy attack and defense framework has been widely investigated, including Zhang et al. (2023a); du Pin Calmon and Fawaz (2012); Zhang et al. (2023b); Kang et al. (2022); Zhang et al. (2023f); Asi et al. (2023).
III-A Privacy Attack Model
The threat model provides a structured perspective on the potential vulnerabilities and risks to client data privacy. Herein, we examine the various facets of an attacker’s strategy, including their intents, resources, intelligence, and tactics.
@@ III-A Privacy Attack Model
Within our framework, we posit a scenario where a protector operates an algorithm $\mathcal{M}(\cdot)$, processing input data $\mathcal{D}$ to yield output $\breve{W}$. This result must then be conveyed to a recipient—who is presupposed to be the attacker. The attacker aspires to extract confidential details regarding $\mathcal{D}$ from the observable $\breve{W}$, while balancing the need to maintain utility within certain bounds. The underlying objective is to refine the attacker’s probabilistic beliefs about $\mathcal{D}$ through analysis of $\breve{W}$.
@@
Our model assumes a semi-honest (or “honest-but-curious”) attacker. Such an adversary adheres to protocol specifications in terms of computations but seeks to glean additional private information by inferring it from $\breve{W}$. By performing statistical inference on the output $\breve{W}$, the attacker aims to ascertain insights into the protector’s confidential data $\mathcal{D}$.
@@
Based on previous descriptions, the attacker has access to the output $\breve{W}$ (though post-transformation by the protector’s mechanism, they receive a modified version $W$). Furthermore, the attacker may possess pre-existing knowledge about the private data $\mathcal{D}$, potentially acquired through past breaches or leaks. Additionally, the attacker is cognizant of the utility constraints associated with their analyses.
@@
With knowledge of a distorted parameter $W$, we assume that the attacker adopts a Bayesian inference attack, i.e., find a data $d$ that optimizes posterior belief, denoted by:
$\arg\max_{d}\text{IH}(d|w)=\arg\max_{d}[\text{I}(w|d)+\text{H}(d)],$ (1)
where $\text{IH}(d|w)$ represents the logarithm of the posterior probability density function $f_{{D}|{W}}(d|w)$, $\text{I}(w|d)$ is the likelihood function’s logarithm based on observed parameter $w$, and $\text{H}(d)$ symbolizes the prior probability distribution’s logarithm $f_{{D}}(d)$. According to Bayes’ theorem, maximizing the logarithm of posterior $f_{D|W}(d|w)$ relative to $D$ involves the aggregation of both $\log(f_{W|D}(w|d))$ and $\log(f_{D}(d))$.
III-B Privacy Protection Model
We now turn our attention to the defensive measures employed by the protector whose design is to shield sensitive datasets from unauthorized exposure.
@@ III-B Privacy Protection Model
To safeguard privacy post-computation by algorithm $\mathcal{M}(\cdot):\mathcal{D}\rightarrow\breve{W}$, the protector needs to devise a privacy-preserving mechanism, which is a stochastic transformation, having the ability to map $\breve{W}\rightarrow W$, where $W$ is a less-informative derivative. This ensures compliance with privacy standards such as $\epsilon$-differential privacy.
@@
It is incumbent upon the protector to transmit $W$ in lieu of $\breve{W}$ to conceal the original data $\mathcal{D}$. To this end, the protector must be adept at generating and communicating a perturbed variant $W$, which duly obscures the private dataset $\mathcal{D}$. This task may be complicated by historical disclosures of $\mathcal{D}$.
@@
The protector is privy to the private dataset $\mathcal{D}$, the computed output $\breve{W}$, and the intricacies of the employed algorithm $\mathcal{M}(\cdot)$. They are also versed in the workings of the privacy-preserving transformation. Additionally, an understanding of the adversary’s utility limitations is crucial for the protector to tailor the distortion of $\breve{W}$ into $W$ effectively, satisfying the informational requirements of the adversary whilst preserving the sanctity of $\mathcal{D}$.
@@
In defence of privacy (i.e., the confidentiality of dataset $\mathcal{D}$), the protector must execute the randomized privacy-preserving mechanism, which transforms $\breve{W}$ to $W$. Execution of the algorithm is expected to equip it with the requisite properties for safeguarding dataset $\mathcal{D}$ against intrusion while adhering to the adopted privacy standard. For instance, optimized algorithms for protecting privacy under the defense and attacking framework have been designed as detailed in the studies by Zhang et al. (2023f) and du Pin Calmon and Fawaz (2012), while Asi et al. (2023) choose to apply exponential mechanism(i.e., add random noise to $\breve{W}$) to protect privacy.
IV Connecting Privacy Measurements and PAC Robustness in Machine Learning
This section lays the theoretical groundwork for understanding the intricate relationships between different privacy notions and their connection to algorithmic robustness. It commences by delving into the theoretical ties among Maximum Bayesian Privacy (MBP), Average Bayesian Privacy (ABP), and Local Differential Privacy (LDP). Through meticulous definitions and theorems, it elucidates how these privacy concepts are interconnected, enabling transitions between various privacy frameworks. Fundamental results establish that, under certain conditions, LDP implies MBP and vice versa. The section also explores the link between MBP and ABP, offering valuable insights for constructing privacy-preserving algorithms. Furthermore, it investigates the relationship between PAC robust learning algorithms and privacy preservation. PAC robust learning, a performance metric ensuring approximately correct learning even with slightly perturbed data, is examined in the context of privacy protection. The section demonstrates how to derive PAC robustness from algorithms satisfying specific privacy standards and how to construct privacy-preserving algorithms from PAC robust ones. These findings provide a solid theoretical foundation for designing robust learning algorithms that simultaneously offer privacy protection capabilities. By bridging the gap between privacy notions and robustness, this section paves the way for developing algorithms that achieve both data confidentiality and resilience to perturbations.
IV-A Metrics for Privacy and Robustness
This section introduces the key metrics used to quantify privacy and robustness in the proposed defense and attack framework. It provides detailed definitions for privacy measures such as Maximum Bayesian Privacy, Average Bayesian Privacy, and Local Differential Privacy, enabling researchers to evaluate the privacy risks associated with different models. The section also defines PAC Robustness as a measure of an algorithm’s consistency against data variations. These metrics form the foundation for developing and assessing algorithms that balance data utility and confidentiality.
IV-A1 Privacy Measurements
Privacy measurement in machine learning has become a crucial area of research, with various approaches proposed to quantify and evaluate the privacy guarantees of different algorithms and models. This section provides a detailed exposition on the metrics used to quantify privacy under the previously discussed privacy attack and defense framework, including concepts such as Maximum Bayesian Privacy, Average Bayesian Privacy, and Local Differential Privacy. These approaches enable researchers to quantitatively evaluate the privacy risks associated with different models and algorithms. As privacy concerns continue to grow, further advancements in privacy measurement techniques are critical for developing privacy-preserving machine learning solutions. For methodologies and analyses concerning the practical estimation of MBP and ABP, please refer to F and G.
First, we define the concept of average Bayesian privacy and explore its correlation with the notion of maximum Bayesian privacy in Section IV-B. These metrics serve as crucial benchmarks in evaluating the resilience of data to inference attacks. The average privacy leakage (Definition IV.1) measures the variation between the adversary’s prior and posterior beliefs on private information. The adversary obtains the posterior belief through mounting the Bayesian inference attack on the protected model information. Thus, we formally call the privacy leakage the Bayesian privacy leakage. The protected model information may lead to the federated model with lower utility and the federated training with less efficiency in certain scenarios.
[…]
We use JS divergence to measure the privacy leakage instead of KL divergence. The advantage of JS divergence over KL divergence is that it is symmetrical, and its square root satisfies the triangle inequality (Endres and Schindelin (2003)). This property facilitates the quantification of the trade-offs. Now we introduce the average Bayesian privacy of the protection mechanism.
Definition IV.1 (Average Bayesian Privacy).
Let $\mathcal{M}$ represent the protection mechanism that maps from the private information $D$ to the model information $W$, and $\mathcal{G}$ represent the attacking mechanism. Let $\epsilon_{p}$ represent the privacy leakage, which is defined as:
$\epsilon_{p}(\mathcal{M},\mathcal{G})=\sqrt{{\text{JS}}(F^{\mathcal{A}}||F^{\mathcal{B}})},$ (2)
where ${\text{JS}}(F^{\mathcal{A}}||F^{\mathcal{B}})=\frac{1}{2}\int_{\mathcal{D}}f^{\mathcal{A}}_{D}(d)\log\frac{f^{\mathcal{A}}_{D}(d)}{f^{\mathcal{M}}_{D}(d)}\textbf{d}\mu(d)+\frac{1}{2}\int_{\mathcal{D}}f^{\mathcal{B}}_{D}(d)\log\frac{f^{\mathcal{B}}_{D}(d)}{f^{\mathcal{M}}_{D}(d)}\textbf{d}\mu(d)$, $F^{\mathcal{A}}$ and $F^{\mathcal{B}}$ represent the attacker’s belief distribution about $D$ upon observing the protected information and without observing any information respectively, and $f_{D}^{\mathcal{M}}(d)=\frac{1}{2}(f^{\mathcal{A}}_{D}(d)+f^{\mathcal{B}}_{D}(d))$. We say $\mathcal{M}$ is $\epsilon$-ABP when $\epsilon_{p}(\mathcal{M},\mathcal{G})\leq\epsilon$.
@@ IV-A1 Privacy Measurements
We operate under the assumption that the private information $D$ takes on discrete values, and then
${\text{JS}}(F^{\mathcal{A}}||F^{\mathcal{B}})$ $=\frac{1}{2}\left[\text{KL}\left(F^{\mathcal{A}}||F^{\mathcal{M}}\right)+\text{KL}\left(F^{\mathcal{B}}||F^{\mathcal{M}}\right)\right]$
$=\frac{1}{2}\sum_{d\in\mathcal{D}}f^{\mathcal{A}}_{D}(d)\log\frac{f^{\mathcal{A}}_{D}(d)}{f^{\mathcal{M}}_{D}(d)}+\frac{1}{2}\sum_{d\in\mathcal{D}}f^{\mathcal{B}}_{D}(d)\log\frac{f^{\mathcal{B}}_{D}(d)}{f^{\mathcal{M}}_{D}(d)}.$
Now we provide the formal definition for $(\epsilon,\beta,\alpha)$-ABP.
Definition IV.2 ($(\epsilon,\beta,\alpha)$-ABP).
Algorithm $\mathcal{M}$ achieves $(\epsilon,\beta,\alpha)$-ABP for a distribution $P$, if it satisfies: (1)being $\epsilon$-ABP, i.e., the privacy leakage $\epsilon_{p,a}\leq\epsilon$, (2)achieving standard accuracy $\alpha$ with a confidence of $1-\beta$.
The Maximum Bayesian Privacy (denoted by $\epsilon_{p,m}$) encapsulates the highest degree of privacy assurance that can be sustained against an adversary with partial or complete knowledge about the data distribution. This metric is situated within the Bayesian privacy paradigm, which assesses the risk associated with the disclosure of private information. It underscores the importance of maintaining privacy when data undergoes sharing or utilization for designated tasks.
Definition IV.3 ($\xi$-Maximum Bayesian Privacy).
Let $F$ denote the belief distribution held by the attacker regarding the private data, and let $f$ indicate the associated probability density function. We define the system to possess $\xi$-Maximum Bayesian Privacy if for any $w\in\mathcal{W}$ and $d\in\mathcal{D}$, the following inequality is satisfied:
$e^{-\xi}\leq\frac{f_{D|W}(d|w)}{f_{D}(d)}\leq e^{\xi}.$
The $\xi$-Maximum Bayesian Privacy quantifies the level of privacy protection afforded by measuring the disparities in the conditional and marginal distributions of the data. A smaller value of $\xi$ implies stronger privacy guarantees against potential adversaries.
Definition IV.4 (PAC Standard Accuracy).
Let $\mathcal{M}$ be an algorithm tasked with estimating a statistic $\mu$, and $d$ represent the private information. The algorithm $\mathcal{M}$ is said to possess standard accuracy $\alpha$ with confidence $1-\beta$ if:
@@
The accuracy is measured within the framework of Probably Approximately Correct (PAC) learning following Zhang et al. (2023c).
Definition IV.5 ($(\epsilon,\beta,\alpha)$-MBP).
Let $\mathcal{M}$ denote an algorithm designed for the estimation of statistic $\mu$. Algorithm $\mathcal{M}$ achieves $(\epsilon,\beta,\alpha)$-MBP for a distribution $P$, if it satisfies two criteria: being $\epsilon$-MBP, and achieving standard accuracy $\alpha$ with a confidence of $1-\beta$. Here, $\alpha$ symbolizes a constant deviation threshold.
IV-A2 PAC Robustness Measurement
In parallel with privacy, we also define robustness within the PAC framework, specifically addressing mean-estimation algorithms.
Definition IV.6 (PAC Robust Accuracy).
@@ IV-A2 PAC Robustness Measurement
Consider $\mathcal{M}$ as an algorithm for estimating statistic $\mu$, and $d$ represent the private information. The algorithm $\mathcal{M}$ is considered to have robust accuracy $\alpha$ with confidence $1-\beta$, if for all datasets $d^{\prime}\neq d$, the following holds with probability at least $1-\beta$:
$\|\mathcal{M}(d^{\prime})-\mu(P)\|\leq\alpha.$ (4)

### v5

@@ 2.5 Relation between Robustness and Privacy
    When corruption ratio $\tau\approx\log(n)/n\epsilon$
@@
    When $\tau\geq\frac{2(d\cdot\log(\frac{R}{\alpha_{0}}+1)+\log\frac{2}{\beta})}{n\epsilon}$
3 Privacy Attack and Defense Framework
This section formalizes the adversarial ecosystem inherent in machine learning data privacy. We systematically establish a comprehensive threat model characterizing adversarial objectives, capabilities, knowledge structures, and methodologies, while concurrently delineating defensive countermeasures for sensitive information protection. This dual-aspect framework builds upon extensive prior research [59, 60, 35, 65].
3.1 Privacy Attack Model
The threat model rigorously conceptualizes vulnerabilities compromising client data privacy through four adversarial dimensions:
@@ 3.1 Privacy Attack Model
Our framework models a protector operating algorithm $\mathcal{M}(\cdot):\mathcal{D}\rightarrow\breve{W}$ and an adversarial recipient aiming to infer confidential aspects of $\mathcal{D}$ from $\breve{W}$ while maintaining functional utility. Formally, the adversary seeks to refine posterior beliefs $\mathbb{P}(\mathcal{D}|\breve{W})$ through analytical scrutiny of the observable output.
@@
We assume a semi-honest ("honest-but-curious") adversary who faithfully executes protocol specifications while illicitly reconstructing $\mathcal{D}$ through statistical analysis of $\breve{W}$, maintaining operational integrity throughout the privacy attack.
@@
The adversary possesses access to the sanitized output $W$ (the transformed version of $\breve{W}$), potential prior knowledge of $\mathcal{D}$ obtained through historical breaches, and understanding of the utility constraints governing analytical operations.
@@
Given observation $W$, the attacker implements Bayesian inference to maximize posterior belief:
$\arg\max_{d}\log f_{D|W}(d|w)=\arg\max_{d}\left[\underbrace{\log f_{W|D}(w|d)}_{\text{log-likelihood}}+\underbrace{\log f_{D}(d)}_{\text{log-prior}}\right]$ (1)
where $f_{D|W}$, $f_{W|D}$, and $f_{D}$ denote the posterior density, likelihood function, and prior distribution respectively.
3.2 Privacy Protection Model
We articulate defensive measures against unauthorized disclosure through four protector dimensions:
@@ 3.2 Privacy Protection Model
The protector engineers a stochastic privacy-preserving mechanism $\mathcal{P}:\breve{W}\rightarrow W$ to generate a carefully degraded information derivative $W$ that satisfies formal privacy guarantees under rigorous information-theoretic metrics.
@@
The protector possesses operational capacity to generate and transmit $W$ instead of $\breve{W}$, obscuring $\mathcal{D}$ through controlled perturbation while contending with potential information leakage from historical disclosures.
@@
The protector maintains complete knowledge of the sensitive dataset $\mathcal{D}$, pristine algorithm output $\breve{W}$, computational semantics of $\mathcal{M}(\cdot)$, operational parameters of $\mathcal{P}(\cdot)$, and crucially, the adversary’s utility constraints.
@@
To safeguard $\mathcal{D}$-confidentiality, the protector deploys a randomized mechanism $\mathcal{P}(\breve{W})$ that produces $W$ satisfying dual objectives:
$\mathbb{E}[\mathcal{U}(W)]$ $\geq\tau$ (Utility preservation)
$\mathcal{I}(\mathcal{D};W)$ $\leq\epsilon$ (Privacy guarantee)
where $\mathcal{U}$ represents the utility function, $\tau$ denotes the minimum utility threshold, $\mathcal{I}$ quantifies privacy leakage, and $\epsilon$ establishes the privacy budget. This framework achieves a provable trade-off between data utility and confidentiality protection, implemented through optimized perturbation algorithms [65].
4 Connecting Privacy Measurements and PAC Robustness in Machine Learning
This section establishes the theoretical foundations connecting privacy frameworks with robustness guarantees in machine learning. We systematically investigate the mathematical relationships between Maximum Bayesian Privacy (MBP), Average Bayesian Privacy (ABP), and Local Differential Privacy (LDP), while demonstrating how these privacy notions fundamentally relate to Probably Approximately Correct (PAC) robustness. Through rigorous definitions and theorems, we prove equivalence relations between privacy frameworks under specific conditions and derive quantitative bounds for converting between privacy metrics. Crucially, we establish that privacy-preserving mechanisms inherently confer robustness properties, providing a unified theoretical framework for designing algorithms that simultaneously ensure data confidentiality and resilience to perturbations. These results bridge the gap between privacy and robustness research communities, offering new principles for developing trustworthy learning systems.
4.1 Measurements for Bayesian Privacy
This subsection formalizes the Bayesian privacy framework for quantifying inference risks in data publishing. We introduce three complementary privacy metrics–Maximum Bayesian Privacy (MBP), Average Bayesian Privacy (ABP), and their relationship to Local Differential Privacy (LDP)–that measure protection against training data reconstruction attacks. These metrics capture different aspects of privacy risk: MBP quantifies worst-case vulnerability, ABP measures expected privacy loss, and LDP provides distributional guarantees. The definitions establish a rigorous mathematical foundation for evaluating privacy-preserving mechanisms, enabling precise comparison between protection techniques and formal analysis of privacy-utility trade-offs.
Definition 4.1 (Average Bayesian Privacy).
Let $\mathcal{M}$ represent the protection mechanism mapping private information $D$ to released information $W$, and $\mathcal{G}$ represent the attacker’s inference mechanism. The privacy leakage $\epsilon_{p,a}$ is quantified using the square root of Jensen-Shannon divergence between:
$\epsilon_{p,a}(\mathcal{M},\mathcal{G})=\sqrt{{\text{JS}}(F^{\mathcal{A}}\parallel F^{\mathcal{B}})}$ (2)
where $F^{\mathcal{A}}$ denotes the attacker’s posterior belief about $D$ after observing $W$, $F^{\mathcal{B}}$ represents the prior belief without observation, and $F^{\mathcal{M}}=\frac{1}{2}(F^{\mathcal{A}}+F^{\mathcal{B}})$ serves as the reference distribution. The mechanism satisfies $\epsilon$-ABP when $\epsilon_{p,a}\leq\epsilon$.
@@ 4.1 Measurements for Bayesian Privacy
For discrete $D$, the JS divergence decomposes as:
${\text{JS}}(F^{\mathcal{A}}\parallel F^{\mathcal{B}})=\frac{1}{2}\left[\text{KL}\left(F^{\mathcal{A}}\parallel F^{\mathcal{M}}\right)+\text{KL}\left(F^{\mathcal{B}}\parallel F^{\mathcal{M}}\right)\right]$
This definition provides a symmetric, bounded measure of privacy leakage that overcomes limitations of KL divergence. By quantifying the information gain between prior and posterior beliefs, it captures the effectiveness of privacy protection against Bayesian inference attacks. The JS divergence formulation ensures the metric is well-defined, symmetric, and satisfies the triangle inequality, enabling rigorous composition and comparison of privacy mechanisms.
Definition 4.2 ($(\epsilon,\beta,\alpha)$-ABP).
A mechanism $\mathcal{M}$ achieves $(\epsilon,\beta,\alpha)$-Average Bayesian Privacy for distribution $P$ if: (1) It satisfies $\epsilon$-ABP ($\epsilon_{p,a}\leq\epsilon$), (2) It maintains standard accuracy $\alpha$ with confidence $1-\beta$.
This definition establishes a comprehensive privacy-accuracy benchmark, requiring mechanisms to simultaneously bound expected privacy leakage while preserving utility. The joint requirement formalizes the dual objectives of privacy-preserving data analysis, where protection must not come at the expense of excessive accuracy degradation.
The following definition formalizes worst-case privacy vulnerability against informed adversaries:
Definition 4.3 ($\xi$-Maximum Bayesian Privacy).
A system provides $\xi$-Maximum Bayesian Privacy if for all $w\in\mathcal{W}$ and $d\in\mathcal{D}$:
$e^{-\xi}\leq\frac{f_{D|W}(d|w)}{f_{D}(d)}\leq e^{\xi}$
This metric quantifies the maximum possible privacy violation by bounding the multiplicative change between prior and posterior beliefs. Smaller $\xi$ values indicate stronger protection against reconstruction attacks, with $\xi=0$ corresponding to perfect privacy where observations reveal no additional information.
Definition 4.4 (PAC Standard Accuracy).
A mechanism $\mathcal{M}$ estimating statistic $\mu$ achieves standard accuracy $\alpha$ with confidence $1-\beta$ if:
@@
This definition formalizes utility preservation within the Probably Approximately Correct framework, requiring estimates to remain close to true values with high probability.
Definition 4.5 ($(\epsilon,\beta,\alpha)$-MBP).
Mechanism $\mathcal{M}$ achieves $(\epsilon,\beta,\alpha)$-Maximum Bayesian Privacy for distribution $P$ if: (1) It satisfies $\epsilon$-MBP per Definition 4.3, (2) It achieves standard accuracy $\alpha$ with confidence $1-\beta$.
This combined criterion establishes a rigorous standard for privacy-preserving estimation, requiring both worst-case privacy protection and utility preservation simultaneously.
4.2 Measurements for PAC Robustness
We now formalize robustness within the PAC framework, complementing the privacy definitions. This robustness metric quantifies algorithmic stability against dataset perturbations, which is crucial for deployment in adversarial environments or non-stationary data distributions.
Definition 4.6 (PAC Robust Accuracy).
@@ 4.2 Measurements for PAC Robustness
A mechanism $\mathcal{M}$ estimating statistic $\mu$ achieves robust accuracy $\alpha$ with confidence $1-\beta$ if for all perturbed datasets $d^{\prime}$:
$\|\mathcal{M}(d^{\prime})-\mu(P)\|\leq\alpha$ (4)
holds with probability at least $1-\beta$ over both mechanism randomness and dataset perturbations. We say $\mathcal{M}$ is $\alpha$-PAC-Robust with confidence $1-\beta$.
This definition requires algorithms to maintain accuracy under arbitrary dataset perturbations, formalizing resilience against data corruption, adversarial examples, and distribution shifts. The probabilistic guarantee ensures reliability even when facing worst-case data manipulations.
4.3 Theoretical Connections between Privacy Notions
We now establish fundamental equivalence relationships between privacy frameworks. These theorems enable translation between privacy metrics and provide theoretical foundations for comparing protection guarantees across different privacy paradigms.
The following lemma establishes the equivalence between Maximum Bayesian Privacy and Local Differential Privacy under bounded prior distributions:
Lemma 4.1 (Generalized Relationship between MBP and LDP adapted from theorem 1 of  [34]).
For any mechanism $\mathcal{M}$:
  • If $\mathcal{M}$ satisfies $\xi$-LDP, then it provides $\xi$-MBP
  • If $\mathcal{M}$ satisfies $\xi$-MBP, then it provides $2\xi$-LDP
This theorem establishes a quantitative equivalence between the Bayesian and differential privacy frameworks. The first implication shows that LDP mechanisms automatically provide Bayesian privacy guarantees, while the converse demonstrates that Bayesian private mechanisms yield differential privacy. The $\epsilon$ parameter quantifies the deviation from uniform priors, with smaller values tightening the equivalence bounds. This result enables direct comparison between privacy frameworks and allows practitioners to leverage results from one paradigm in the other.
The next theorem quantifies how worst-case privacy protection (MBP) bounds expected privacy leakage (ABP):
Theorem 4.2 (Generalized Relationship Between MBP and ABP).
Assume the attacker’s prior belief satisfies $\frac{f^{\mathcal{B}}_{D}(d)}{f_{D}(d)}\in[e^{-\epsilon},e^{\epsilon}]$. Any $\epsilon_{p,m}$-MBP mechanism bounds the ABP leakage by:
$\epsilon_{p,a}\leq\frac{1}{\sqrt{2}}\sqrt{(\epsilon_{p,m}+\epsilon)\left(e^{\epsilon_{p,m}+\epsilon}-1\right)}$
This result establishes that worst-case privacy protection (MBP) implies bounds on average-case privacy leakage (ABP). The bound tightens as $\epsilon_{p,m}$ decreases, showing that strong MBP guarantees necessarily enforce strong ABP protection. The $\epsilon$ parameter captures prior distribution mismatch, with exact priors ($\epsilon=0$) yielding the tightest bound. This theorem provides crucial guidance for mechanism design: by optimizing for MBP, designers automatically obtain ABP guarantees, simplifying the privacy certification process.
4.4 PAC Robustness of Bayesian Private Algorithms
We now establish the foundational connection between privacy protection and algorithmic robustness. The following theorem demonstrates that privacy mechanisms inherently confer robustness properties:
Theorem 4.3 (PAC Robustness of $\epsilon$-MBP Protection Mechanisms).